My shower is leaking at a rate of one drop per second.
A. how many cups of water per hour are being wasted?
B. how many gallons per week?
useful conversion; 1gallon = 4 quarts, 4 cups = 1 quart,
1 cup = 16tbsp, 1tbsp = 3tsp,
1tsp = 16drops.

Found 2 solutions by Mathtut, laflorida:
Answer by Mathtut(3670)

1gallon = 4 quarts, 4 cups = 1 quart, 1 cup = 16tbsp, 1tbsp = 3tsp, 1tsp = 16drops.
:
tsp=16drops and there is 60 drops per minute...so 60/16=3.75 tsp/min....that is where we will begin as there is 1 drop per second.
:
A
:
B

Answer by laflorida(4)

My shower is leaking at a rate of one drop per second.
A. how many cups of water per hour are being wasted?
B. how many gallons per week?
useful conversion; 1gallon = 4 quarts, 4 cups = 1 quart, 1 cup = 16tbsp, 1tbsp = 3tsp, 1tsp = 16drops.
60 mins x 60 seconds = 3600 seconds per hour
3600 drops per hour = 3600/16 = 225 tsp
225/3 = 75 tbsp
75/16 = 4.6875 cups Per Hour
4.6875/4 = 1.171875 quarts
1.171875/4 = 0.29296875 Gallons Per Hour
24 hrs per day x 7 days in a week = 168 hrs per week
0.29296875 Gallons Per Hour x 168 hrs = 49.21875
A. 4.6875 cups Per Hour
B. 49.21875 Gallons per Week